OpenStudy (anonymous):

If f(x) = 3x + 5, then what is f-1(x) ?

OpenStudy (abb0t):

Do you mean \(f^{-1}(x)\)

geerky42 (geerky42):

To find inverse function, switch x and y to each other in original function then isolate y.

geerky42 (geerky42):

So \(f(x) = y = 3x + 5\) right? So \(f^{-1}(x) \Rightarrow x = 3y + 5\) \(x - 5 = 3y\) \(f^{-1} = y = \boxed{\dfrac{1}{3}x - \dfrac{5}{3}}\)

geerky42 (geerky42):

Is this clear? @stevensde

OpenStudy (anonymous):

Yes, thank you very much.

geerky42 (geerky42):

You're welcome.

OpenStudy (abb0t):

It's not \(that\) clear..

geerky42 (geerky42):

Not for you.